OpenStudy (anonymous):

Two trains left the station at the same time, one heading due east at 40 miles per hour and the other heading due west at 50 miles per hour. How far apart were they after 4 hours? A. 160 B. 200 C. 360 D. 400

OpenStudy (anonymous):

well you have to see how far they have gotten after 4 hours. so 4*40=160 50*4=200. since it was opposite directions you add. so 160+200=360.